OpenStudy (thebadyboy):

During his first year of life, Avery gained 12 3/4 lb. During his second year of life, Avery gained 40% less weight than he did during his first year of life. How much weight did Avery gain during his second year of life? Express your answer as a mixed number in simplest form.

OpenStudy (thebadyboy):

\[7\frac{ 13 }{ 20}\]

OpenStudy (thebadyboy):

right?

OpenStudy (thebadyboy):

@StudyGurl14 @surjithayer

OpenStudy (domebotnos):

i love your picture ;)

OpenStudy (thebadyboy):

thx

OpenStudy (thebadyboy):

so yes

OpenStudy (studygurl14):

Sorry, that's incorrect. Do this: \(\large 40\%\times 12\frac{3}{4}\) (hint: convert 40% into a decimal and \(12\frac{3}{4}\) into an improper fraction)

OpenStudy (thebadyboy):

\[5\frac{ 1 }{ 10}\]

OpenStudy (thebadyboy):

@surjithayer

OpenStudy (thebadyboy):

yes or no

OpenStudy (thebadyboy):

r u there

OpenStudy (anonymous):

you are correct it is 60% of 12 3/4

OpenStudy (thebadyboy):

thx

OpenStudy (anonymous):

yw